OpenStudy (anonymous):

for the given function, compute the average rate of change f(x2)-f(x1)/x2-x1 with x1= a and x2 = a+h 1) f(x) = 7-2x

OpenStudy (anonymous):

\(f\) is a line, and the average rate of change of a line is a constant, it is the slope

OpenStudy (anonymous):

what you are being asked for is the slope of \(y=7-2x\)

OpenStudy (anonymous):

how do i compute that in the x1 and x2 equations? I dont really get this concept

OpenStudy (anonymous):

the slope of \(y=7-2x\) is \(-2\)

OpenStudy (anonymous):

you are supposed to write \[\frac{f(a+h)-f(a)}{a+h-a}\] or\[\frac{7-2(a+h)-(7-2a)}{h}\]

OpenStudy (anonymous):

when you clean up the numerator after some algebra, you will end up with \[\frac{-2h}{h}=-2\]

OpenStudy (anonymous):

ok, thank you!